OpenStudy (anonymous):

Find the general solution of the following:

OpenStudy (anonymous):

\[\sin (x+ 40) = 1/\sqrt{2}\] \[\sin5x = \sin 3x\] \[\tan ax = \cot bx \] \[\tan ^{2}3x = 3\]

OpenStudy (anonymous):

\[x+40 {}^{\circ}=\sin ^{-1}\left(\frac{1}{\sqrt{2}}\right) \]\[x=\sin ^{-1}\left(\frac{1}{\sqrt{2}}\right)-40 {}^{\circ} \]\[x=\frac{\pi }{4}-\frac{\pi (40 {}^{\circ})}{180 {}^{\circ}} \]\[x=\frac{\pi }{36} \]

OpenStudy (anonymous):

tan 3x = \[\sqrt{3}\] , 3x = \[\tan^{-1} \sqrt{3}\] x=\[ \pi/9\]